OpenStudy (anonymous):

The sum of -5, -9, and 10 is

Nnesha (nnesha):

negative 5 and negative 9 = ??

Nnesha (nnesha):

sum mean add

OpenStudy (anonymous):

Oh -4

Nnesha (nnesha):

-5 and -9 is not 4 negative and negative become positive so add -5 and -9

Nnesha (nnesha):

-5-9 =??

OpenStudy (anonymous):

4

OpenStudy (gorv):

10-9=?

Nnesha (nnesha):

-5 and -9 is = -14 so 10 + (-14) = 4 yep you right final answer is 4